The machining division of ITA International has a capacity of 2,010 units. Its sales and cost data are:

Selling price per unit $ 80
Variable manufacturing costs per unit 30
Variable selling costs per unit 3
Total fixed manufacturing overhead 216,600



The machining division is currently selling 1,810 units to outside customers, and the assembly division of ITA International wants to purchase 400 units from machining. If the transaction takes place, the variable selling costs per unit on the units transferred to assembly will be $0/unit, and not $3/unit. What should be the transfer price in order not to affect the machining division’s current profit? (Round answer to 2 decimal places e.g. 5.25.)

Minimum transfer price $

Entry field with incorrect answer now contains modified data
Answer #1

Free capacity available = 2010 -1810 = 200 units

Total minimum transfer price = selling price to outside customer for 200 units - selling cost on 200 units + variable manufacturing cost for 200 units

= 200*80 - 200*3 + 200*30

= $21400

Minimum transfer price = 21400/400 = $53.5 per units